How do I find the point where the graph intersects with the x-axis for the problem: f(x) = √(t+4) - 1
Thank you

Answer by amarjeeth123(569)

f(x) =y= √(t+4) - 1
Where it intersects the x axis y=0
√(t+4)-1=0
√(t+4)=1
t+4=1
t=-3
It intersects the x-axis at t=-3.